OpenStudy (aravindg):

Prove that cos(2pi/15)*cos(4pi/15)*cos(8pi/15)*cos(16pi/15)=1/16

OpenStudy (anonymous):

it's true ! but idk how to prove

OpenStudy (anonymous):

because i have no idea how to write it on here

OpenStudy (anonymous):

If you know what each of the cos(xπ/15) equals you could write those down being multiplied and then do out the work

OpenStudy (phi):

sin(a+b)= sin(a)cos(b)+sin(b)cos(a) (1) for a=b, sin(2a)= 2 sin(a) cos(a) (2) let a= 16 pi/15 (3) (so 2a = 32 pi/15 ) then using (3) in (2), we have sin(2a)= 2 sin(a) cos(a) = 2 (2 sin(a/2) cos(a/2)) cos(a) = 2 (2 (2 sin(a/4) cos(a/4)) cos(a/2)) cos(a) = 2 (2 (2 (2 sin(a/8) cos(a/8)) cos(a/4)) cos(a/2)) cos(a) = 16 sin(a/8) ( cos(a/8) cos(a/4) cos(a/2) cos(a)) now note sin(2a) = sin(2 pi/15) and sin(a/8)= sin(2 pi/15) so, cos(a/8) cos(a/4) cos(a/2) cos(a)= 1/16 or, replacing a with 16 pi/15, cos(2pi/15)*cos(4pi/15)*cos(8pi/15)*cos(16pi/15)=1/16